OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $9.60. If the number of nickels was tripled and the number of dimes was decreased by 31, the value of the coins would be $6.70. How many nickels and dimes does he have?

OpenStudy (anonymous):

write a system of equations; the first one about the value of the money and the second about the number of coins. .05x + .10y = 9.60 and 3x + y-31=6.70

OpenStudy (anonymous):

ok thank you so much I forgot the money... i really need to finish my final, but also need to go to bed unfortunately the two don't mix well.

OpenStudy (anonymous):

Do you have any questions?

OpenStudy (anonymous):

.05n+.10d=9.60 -.05(3n+d-31=6.70) 3(.05n+.10d=9.60) -.15n-.05d+1.55=-.335 .15n+.3d=28.8 .15n-.05d+1.55+.15n+.3d=-.335+28.8 .25d+1.55=28.465 .25d=26.915 d=107.66 3n+107.66-31=6.70 3n+76.66=6.70 3n=-69.96 n=-23.32 is this correct?

OpenStudy (anonymous):

This can't be correct because the answers must be whole numbers and you cannot have a negative amount of coins...

OpenStudy (anonymous):

I know what am I doing wrong?

OpenStudy (anonymous):

give me a minute to work it out....

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

ok. this is what I got: .05n + .10d=9.60 and .05(3n)+.10(d-31)=6.70 then I solved the second eq in terms of d and got d=.98-1.5n Then I substituted in the first and solved. .05n + .10 (.98-1.5n)=9.60 -.10n=-.20 n= 2 nickels; then plug it back in the first one and I got 95 dimes...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yay!